OpenStudy (anonymous):

Help ASAP, MEDAL GIVING Mei's salary starts at 16000 per year with annual raises of 1500. Janets starting salary is 19300 with annual raises of 950. After how many years will the two women be earning the same salary.

OpenStudy (kkutie7):

Variables: m = Mei's earnings j = Janet's earnings y = number of years Given: Mei's salary starts at $16,000 per year with annual raises of $1,500 Means: m = 16000 + 1500y Given: Janet's starting salary is $19,300 with annual raises of $950 Means: j = 19300 + 950y

OpenStudy (kkutie7):

Given: After how many years will the two women be earning the same salary? Means: you need to find out when m = j You have 3 equations. m = 16000 + 1500y j = 19300 + 950y m = j

OpenStudy (kkutie7):

Substitute m with 16000 + 1500y and j with 19300 + 950y in the third equation. m = j 16000 + 1500y = 19300 + 950y Solve. 1500y - 950y = 19300 - 16000 550y = 3300 y = 3300 / 550 y = 6 ANSWER: They will earn the same after 6 years.